Starfax, Inc., manufactures a small part that is widely used in various electronic products such as home computers. Results for the first three years of operations were as follows (absorption costing basis):

Year 1 Year 2 Year 3
Sales $ 1,100,000 $ 850,000 $ 1,100,000
Cost of goods sold 850,000 600,000 900,000
Gross margin 250,000 250,000 200,000
Selling and administrative expenses 220,000 190,000 220,000
Net operating income (loss) $ 30,000 $ 60,000 $ (20,000 )

  

In the latter part of Year 2, a competitor went out of business and in the process dumped a large number of units on the market. As a result, Starfax’s sales dropped by 20% during Year 2 even though production increased during the year. Management had expected sales to remain constant at 50,000 units; the increased production was designed to provide the company with a buffer of protection against unexpected spurts in demand. By the start of Year 3, management could see that it had excess inventory and that spurts in demand were unlikely. To reduce the excessive inventories, Starfax cut back production during Year 3, as shown below:

Year 1 Year 2 Year 3
Production in units 50,000 60,000 40,000
Sales in units 50,000 40,000 50,000

Additional information about the company follows:

  1. The company’s plant is highly automated. Variable manufacturing expenses (direct materials, direct labor, and variable manufacturing overhead) total only $5.00 per unit, and fixed manufacturing overhead expenses total $600,000 per year.

  2. A new fixed manufacturing overhead rate is computed each year based that year's actual fixed manufacturing overhead costs divided by the actual number of units produced.

  3. Variable selling and administrative expenses were $3 per unit sold in each year. Fixed selling and administrative expenses totaled $70,000 per year.

  4. The company uses a FIFO inventory flow assumption. (FIFO means first-in first-out. In other words, it assumes that the oldest units in inventory are sold first.)

Starfax’s management can’t understand why profits doubled during Year 2 when sales dropped by 20% and why a loss was incurred during Year 3 when sales recovered to previous levels.

Required:

1. Prepare a contribution format variable costing income statement for each year.

2. Refer to the absorption costing income statements above.

a. Compute the unit product cost in each year under absorption costing. Show how much of this cost is variable and how much is fixed.

b. Reconcile the variable costing and absorption costing net operating income figures for each year.

5b. If Lean Production had been used during Year 2 and Year 3, what would the company’s net operating income (or loss) have been in each year under absorption costing?
Answer #1

Part 1

Year 1

Year 2

Year 3

Sales

1100000

850000

1100000

Variable expenses:

Variable cost of goods sold (50000*5=250000); (40000*5=200000); (50000*5=250000)

250000

200000

250000

Variable selling and administrative expenses (50000*3=150000); (40000*3=120000); (50000*3=150000)

150000

120000

150000

Total variable expenses

400000

320000

400000

Contribution margin

700000

530000

700000

Fixed expenses:

Fixed manufacturing overhead

600000

600000

600000

Fixed selling and administrative expenses

70000

70000

70000

Total fixed expenses

670000

670000

670000

Net operating income (loss)

$30000

$(140000)

$30000

Part 2 A

Year 1

Year 2

Year 3

Variable manufacturing costs

5

5

5

Fixed manufacturing costs

12 (600000/50000)

10 (600000/60000)

15 (600000/40000)

Unit product cost

$17

$15

$20

Part 2 B

Year 1

Year 2

Year 3

Variable costing net operating income (loss)

30000

(140000)

30000

Add (Deduct): Fixed manufacturing overhead cost deferred in inventory from Year 2 to Year 3 under absorption costing (10*(60000-40000))

200000

(200000)

Add: Fixed manufacturing overhead cost deferred in inventory from Year 3 to the future under absorption costing (10000*15) (40000+20000-50000)

150000

Absorption costing net operating income (loss)

30000

60000

(20000)

Part 5 B

Year 1

Year 2

Year 3

Unit sales

50000

40000

50000

Sales

1100000

850000

1100000

Cost of goods sold

Cost of goods manufactured

850000

680000

850000

Add underapplied overhead

120000

Cost of goods sold

850000

800000

850000

Gross margin

250000

50000

250000

Selling and administrative expenses

220000

190000

220000

Net operating income (loss)

30000

(140000)

30000

40000*17 = 680000

(50000-40000)*12 = 120000